# The anisotropic expansion rate of the local Universe and its covariant cosmographic interpretation

### v1

@@ 1 Introduction
This work is the fifth contribution in an ongoing series [42, 43, 44, 45], through which we aim to open a new window on studies of the local expansion rate by developing a new method to characterize it in a model-independent and non-perturbative manner, without suming the FLRW metric, Einstein’s field equations, or invoking concepts such as peculiar velocities or density fluctuations. Central to this framework is the expansion rate fluctuation field, $\eta$ (introduced in [42, 44]), a scalar Gaussian observable designed to identify and classify deviations from isotropy in the redshift–distance relation. An important caveat is in order. With this cosmological observable, we are not probing perturbations of the Hubble constant $H_{0}$, which would require assuming a fixed background and relying on FLRW modeling in the estimation of expansion-rate anisotropies. Although our fully model-independent approach does not provide a specific value of $H_{0}$, its key advantage is that it remains unaffected by distance-dependent selection biases, such as the Malmquist bias.
@@
The purpose of this paper is threefold.
@@ 2.1 Cosmicflows-4 sample
Figure 1 illustrates the redshift dependence of morphological composition, with early-type galaxies becoming increasingly dominant at larger redshifts. The angular distribution of objects (top panel of Figure 2 shows an isotropic coverage within $z\sim 0.03$, moderately uniform sampling out to $z\sim 0.05$, and strong anisotropies beyond this scale and up to $z\sim 0.1$. The Galactic plane blocks observations within $\sim 10^{\circ}$, producing the so-called Zone of Avoidance (ZoA)
@@ 3 Expansion rate fluctuation field in the local Universe
By design, $\eta$ is thus a random Gaussian fluctuating variable with zero mean. It attains non-vanishing values if cosmic expansion is not perfectly isotropic and homogeneous. Small departures from zero are expected due to local structure. Significant departures can indicate breakpoints of the Cosmological Principle, specially if these happen at higher redshift bins. The method for estimating the expansion rate fluctuation field at a galaxy’s position, as well as the optimal choice of redshift shell thickness to ensure that the quantity is monopole-free, are discussed in detail in Appendix A.
@@
The two-dimensional expansion rate fluctuation field traced by each sample exhibits significant anisotropy, with fluctuations displaying coherent large-scale angular correlations rather than random noise. In the lowest redshift bin (CF4a), and for the given resolution of the reconstruction (see second row in Figure 2, the expansion rate fluctuation peaks at $\tilde{\eta}\sim 0.05\pm 0.01$ in a direction near the Galactic plane, indicating that, at fixed distances, redshifts are systematically higher than average in that region. If interpreted within the framework of the standard cosmological model, this would imply a local variation in the Hubble constant, relative to a fiducial value of $H_{0}=70$km/s/Mpc, at the level of approximately $10\%$, i.e. $7$ km/s/Mpc. Notably, a similar anisotropic pattern is qualitatively preserved across increasing redshift shells. However, in the highest redshift bin (CF4d), the relevant region of maximal expansion indicated by the lower-redshift shells is not sampled due to the limited sky coverage of the CF4d sample.
@@ 4.1 Robustness of the results
Interestingly, the Pantheon+ sample, the sparsest of all those shown in Figure 3, also exhibits the same dipolar and quadrupolar patterns identified in the galaxy samples. However, its quadrupole intensity is very weak, becoming apparent only when guided by the structures already revealed in the denser datasets. Since all the Pantheon+ objects are also included in the CF4SN sample (indeed they constitute a large fraction of this sample, around $63\%$ at $0.01<\tilde{z}<0.05$). This raises the question of whether the difference in intensity of the quadrupole observed in the CF4SN and Pantheon+ is statistical in nature, resulting solely from the different sample sizes.
@@ 5.1 Bulk motion of matter
One crucial point to highlight is that, if the Shapley supercluster is the dominant structure in the region, the multipoles in the deepest redshift shell ($0.05<\tilde{z}<0.1$) should reverse their polarity and their amplitude should decay [44]. However, Figures 2 and 6 show that the structure of the dipole and quadrupole components (both in amplitude and orientation) remain consistent in the deepest redshift shells, which survey the region beyond the Shapley supercluster. This provides strong evidence for the existence of a coherent, large-scale structure beyond $z\approx 0.05$, within which the local volume is embedded.
@@
Given the tensions that arise when interpreting our findings within the framework of Standard Model predictions, we now adopt a fully model-independent approach to analyze the observed fluctuations in the expansion rate field and to extract their physical meaning: the covariant cosmographic framework.
@@ 6 Interpreting the anisotropies in the framework of covariant cosmography
An important—though often overlooked (see, e.g., [72])—aspect in theoretical treatments of the reliability of the CC formalism is the practical distinction between its formal and operational definitions. While the higher-order CC parameters are formally defined as successive derivatives of the luminosity distance with respect to redshift (in the matter frame) evaluated at the observer’s position, their actual estimation relies not on derivatives at the observer, but on fitting. Specifically, the luminosity distance is treated as an explicit function of redshift over a finite interval, and the CC parameters are obtained as the best-fit coefficients that minimize the discrepancy between the model and observed distance-redshift data within that interval (see also [45]). While these parameters are defined locally, they are estimated non-locally to reduce sensitivity to local noise in the surroundings of the observer and to be more sensitive to the large-scale structure of the gravitational field. In this sense, for example, although the Hubble constant $H_{0}$ is formally defined as the first-order derivative of redshift with respect to distance at the observer’s position in the standard model, it is empirically determined as the slope of the distance–redshift relation over a finite redshift interval. This technique offers three main advantages. Theoretically, it avoids reliance on model-dependent estimates of matter density fields or on the ill-defined problem of averaging over an a priori unknown background spacetime. Observationally, it yields more stable estimates across a broader redshift range, with reliability that can be verified a posteriori through goodness-of-fit tests or consistency checks. Moreover this approach also provides an estimate of the minimum radius around the observer within which data must be excluded from the analysis, if one is to recover the large scale geometry of space. Physically, this means that one can quantify directly from data the characteristic scale at which the CC formalism breaks down due to the failure of the fluid limit approximation on which it relies.
@@ 6.1 Results
In contrast, the jerk dipole $\mathbb{J}_{2}$ is poorly constrained in all three samples, with uncertainties exceeding the central values. As a result, no definitive conclusions can be drawn about third-order anisotropies in the expansion from these measurements alone. Also the structural combinations $\mathbb{S}_{1}+4\mathbb{Q}_{1}(2\mathbb{J}_{0}-\mathbb{R}_{0})$ and $\mathbb{S}_{3}+4\mathbb{Q}_{3}(2\mathbb{J}_{0}-\mathbb{R}_{0})$, which encode nonlinear couplings between dipole and octupole terms, are moderately significant in CF4 but suffer from large uncertainties in the CF4TF and CF4FP subsamples. While nonlinear structure is likely present in the full dataset, it is difficult to isolate with higher precision in the individual subsets.
@@ 6.2 Robustness of the results
Figure 14: The best-fit covariant cosmographic parameters and the observer’s velocity $v_{o}$ as a function of the lower redshift cut ($\tilde{z}_{\min}$) for the CF4 sample. The black lines presents the best values chosen for Table 4 and correspond to those estimated using $\tilde{z}_{\min}=0.01$. The shaded area indicates the range excluded to estimate the parameters quoted in Table 4.
@@
6.3 Modelling cosmic matter as a dust fluid
@@
Including very local data ($\tilde{z}<0.01$) compromises the ability of the covariant cosmographic model to describe expansion rate fluctuations across the entire volume considered ($\tilde{z}<0.1$). The fundamental physical reason for this lies in the breakdown of the continuum approximation on which the formalism is based. The CC framework assumes that the cosmic matter field can be modeled as a smooth pressure-free fluid (‘dust’) on sufficiently large scales. On smaller scales however, the data are dominated by the ‘thermal’ motions (i.e. velocity dispersion) of galaxies, which reflect local gravitational fluctuations rather than the large-scale cosmic gravity. This not only prevents the calculation of CC parameters via local derivatives, but also highlights a further limitation: the formalism does not provide any a priori method for determining the scale below which the fluid approximation fails.
Remarkably, as shown in Figure 14, this breakdown scale can be identified a posteriori using the data (see also Figures 11 and 15). Observationally, this transition consistently occurs around $\tilde{z}\sim 0.01$ across all modes in which the expansion rate fluctuation field is decomposed: its dipole, quadrupole, and octupole all begin to diverge coherently from theoretical expectations below this redshift. This finding serves as a caution against defining the matter comoving observer in the covariant cosmography framework by averaging data within spherical volumes centered on the Local Group, with radii of approximately $30\,h^{-1}\,\mathrm{Mpc}$ (where, for example, the CF4 sample typically includes nearly 3000 objects.)
@@
A more physically motivated estimate of the spatial scale over which one must average to define the velocity of the matter element representing the CC observer can be inferred through a model-dependent argument based on standard cosmological information. We know that the velocity of the matter observer is $v_{o}=188\pm 22$ km/s in the direction $(l,b)=(299^{\circ},5^{\circ})$. Using Cosmicflows-3 sample, [74] determined the contributions to the Local Group (LG) velocity from mass enclosed within a sphere of radius $38$ Mpc, and from the surrounding shell extending from $38$ to $100$ Mpc (see Table 5). Given the observed CMB dipole, we can deduce that the component of the LG’s motion generated by mass beyond $38$ Mpc is approximately $v_{38}=320\pm 70$ km/s in the direction $(l,b)=(292^{\circ},4^{\circ})$, while that generated by structures beyond $100$ Mpc is about $v_{100}=100\pm 70$ km/s in the direction $(l,b)=(302^{\circ},-38^{\circ})$. The fairly good alignment between these directions and the measured $v_{o}$, and the fact that $v_{100}<v_{o}<v_{38}$ suggests that the effective averaging length corresponding to the size of the fluid element representing the CC observer lies somewhere in the range $38<R~(\mathrm{Mpc})<100$. This holds under the simplifying assumption that the contribution of shells external to the volume $V$ provides, within statistical noise, a fair representation of the bulk velocity of $V,$ rather than merely the velocity of the Local Group.

### v2

@@ 1 Introduction
This work is the fifth contribution in an ongoing series [42, 43, 44, 45], through which we aim to open a new window on studies of the local expansion rate by developing a new method to characterize it in a model-independent and non-perturbative manner, without assuming the FLRW metric, Einstein’s field equations, or invoking concepts such as peculiar velocities or density fluctuations. Central to this framework is the expansion rate fluctuation field, $\eta$ (introduced in [42, 44]), a scalar Gaussian observable designed to identify and classify deviations from isotropy in the redshift–distance relation. An important caveat is in order. With this cosmological observable, we are not probing perturbations of the Hubble constant $H_{0}$, which would require assuming a fixed background and relying on FLRW modeling in the estimation of expansion-rate anisotropies. Although our fully model-independent approach does not provide a specific value of $H_{0}$, its key advantage is that it remains unaffected by distance-dependent selection biases, such as the Malmquist bias.
@@
The purpose of this paper is twofold.
@@ 2.1 Cosmicflows-4 sample
Figure 1 illustrates the redshift dependence of morphological composition, with early-type galaxies becoming increasingly dominant at larger redshifts. The angular distribution of objects (top panel of Figure 2 shows an isotropic coverage within $\tilde{z}\sim 0.03$, moderately uniform sampling out to $\tilde{z}\sim 0.05$, and strong anisotropies beyond this scale and up to $\tilde{z}\sim 0.1$. The Galactic plane blocks observations within $\sim 10^{\circ}$, producing the so-called Zone of Avoidance (ZoA)
@@ 3 Expansion rate fluctuation field in the local Universe
By design, $\eta$ is thus a random fluctuating variable with zero mean. It attains non-vanishing values if cosmic expansion is not perfectly isotropic and homogeneous. Small departures from zero are expected due to local structure. Significant departures can indicate breakpoints of the Cosmological Principle, specially if these happen at higher redshift bins. Statistically, $\eta$ follows the same distribution as the distance moduli. In the remainder of this paper, we assume the distance moduli to be Gaussian distributed. This is a practical approximation that works well for high signal-to-noise photometric data and for modelling the scatter associated with calibrating the distance-indicator zero point; see however [56] for a discussion of possible subtle deviations from Gaussianity.
@@
The two-dimensional expansion rate fluctuation field traced by each sample exhibits significant anisotropy, with fluctuations displaying coherent large-scale angular correlations rather than random noise. In the lowest redshift bin (CF4a), and for the given resolution of the reconstruction (see second row in Figure 2), the expansion rate fluctuation peaks at $\tilde{\eta}\sim 0.05\pm 0.01$ in a direction near the Galactic plane, indicating that, at fixed distances, redshifts are systematically higher than average in that region. If interpreted within the framework of the standard cosmological model, this would imply a local variation in the Hubble constant, relative to a fiducial value of $H_{0}=70$km/s/Mpc, at the level of approximately $10\%$, i.e. $7$ km/s/Mpc. Notably, a similar anisotropic pattern is qualitatively preserved across increasing redshift shells. However, in the highest redshift bin (CF4d), the relevant region of maximal expansion indicated by the lower-redshift shells is not sampled due to the limited sky coverage of the CF4d sample.
@@ 4.1 Robustness of the results
Interestingly, the Pantheon+ sample, the sparsest of all those shown in Figure 3, also exhibits the same dipolar and quadrupolar patterns identified in the galaxy samples. However, its quadrupole intensity is very weak, becoming apparent only when guided by the structures already revealed in the denser datasets. All the Pantheon+ objects are also included in the CF4SN sample – indeed they constitute a large fraction of this sample, around $63\%$ at $0.01<\tilde{z}<0.05$. This raises the question of whether the difference in intensity of the quadrupole observed in the CF4SN and Pantheon+ is statistical in nature, resulting solely from the different sample sizes.
@@ 5.1 Bulk motion of matter
One crucial point to highlight is that, if the Shapley supercluster is the dominant structure in the region, the multipoles in the deepest redshift shell ($0.05<\tilde{z}<0.1$) should reverse their polarity and their amplitude should decay [44, 66]. However, Figures 2 and 6 show that the structure of the dipole and quadrupole components (both in amplitude and orientation) remain consistent in the deepest redshift shells, which survey the region beyond the Shapley supercluster. This provides strong evidence for the existence of a coherent, large-scale structure beyond $z\approx 0.05$, within which the local volume is embedded.
@@
A more in-depth analysis is required to stress-test the robustness of the standard-model scenario—for example, a comprehensive comparison of the quadrupole and higher-order moments with Standard Model predictions. Such an analysis falls outside the scope of this paper and is addressed in a separate work [67]. However, given the hints of a possible tension that emerge when our findings are interpreted within the Standard Model framework, we now adopt a fully model-independent approach to analyze the observed fluctuations in the expansion rate field and to extract their physical meaning: the covariant cosmographic (CC) framework.
@@ 6 Interpreting the anisotropies in the framework of covariant cosmography
An important aspect in theoretical treatments of the reliability of the CC formalism is the practical distinction between its formal and operational definitions (see e.g., [75]). While the higher-order CC parameters are formally defined as successive derivatives of the luminosity distance with respect to redshift (in the matter frame) evaluated at the observer’s position, their actual estimation relies not on derivatives at the observer, but on fitting. Specifically, the luminosity distance is treated as an explicit function of redshift over a finite interval, and the CC parameters are obtained as the best-fit coefficients that minimize the discrepancy between the model and observed distance-redshift data within that interval (see also [45]). While these parameters are defined locally, they are estimated non-locally to reduce sensitivity to local noise in the surroundings of the observer and to be more sensitive to the large-scale structure of the gravitational field. In this sense, for example, although the Hubble constant $H_{0}$ is formally defined as the first-order derivative of redshift with respect to distance at the observer’s position in the standard model, it is empirically determined as the slope of the distance–redshift relation over a finite redshift interval. This technique offers three main advantages. Theoretically, it avoids reliance on model-dependent estimates of matter density fields or on the ill-defined problem of averaging over an a priori unknown background spacetime. Observationally, it yields more stable estimates across a broader redshift range, with reliability that can be verified a posteriori through goodness-of-fit tests or consistency checks. Moreover this approach also provides an estimate of the minimum radius around the observer within which data must be excluded from the analysis, if one is to recover the large scale geometry of space. Physically, this means that one can quantify directly from data the characteristic scale at which the CC formalism breaks down due to the failure of the fluid limit approximation on which it relies.
@@ 6.1 Results
In contrast, the jerk quadrupole $\mathbb{J}_{2}$ is poorly constrained in all three samples, with uncertainties exceeding the central values. As a result, no definitive conclusions can be drawn about third-order anisotropies in the expansion from these measurements alone. Also the structural combinations $\mathbb{S}_{1}+4\mathbb{Q}_{1}(2\mathbb{J}_{0}-\mathbb{R}_{0})$ and $\mathbb{S}_{3}+4\mathbb{Q}_{3}(2\mathbb{J}_{0}-\mathbb{R}_{0})$, which encode nonlinear couplings between dipole and octupole terms, are moderately significant in CF4 but suffer from large uncertainties in the CF4TF and CF4FP subsamples. While nonlinear structure is likely present in the full dataset, it is difficult to isolate with higher precision in the individual subsets.
@@ 6.2 Robustness of the results
Figure 14: The best-fit covariant cosmographic parameters and the observer’s velocity $v_{o}$ as a function of the lower redshift cut ($\tilde{z}_{\min}$) for the CF4 sample. The black lines presents the best values chosen for Table 4 and correspond to those estimated using $\tilde{z}_{\min}=0.01$. The shaded area indicates the range excluded to estimate the parameters quoted in Table 4. Note that the multipoles of the CC parameters are constants unlike those of the expansion rate fluctuation field shown in Figure 11.
@@
6.3 Modelling cosmic matter in the CC framework
@@
Including very local data ($\tilde{z}<0.01$) compromises the ability of the covariant cosmographic model to describe expansion rate fluctuations across the entire volume considered ($\tilde{z}<0.1$). The fundamental physical reason for this lies in the breakdown of the smooth continuum approximation on which the formalism is based. The CC framework assumes that the cosmic matter field can be modeled as a smooth pressure-free fluid (‘dust’) on sufficiently large scales. On smaller scales, however, the data are dominated by the ‘thermal’ motions (i.e. velocity dispersion) of galaxies, which reflect local gravitational fluctuations rather than the large-scale cosmic gravity. This not only prevents the calculation of CC parameters using local derivatives but also reveals another limitation: the formalism does not provide any a priori way to determine, in advance, the scale below which the CC framework breaks down. In other words, there is no way to determine in advance the smallest scale at which the anisotropies in the expansion rate no longer accurately represent those of the cosmic volume under study. This issue is similar to a limitation of the standard cosmological model: the assumption of the Cosmological Principle does not specify the scale below which the matter density ceases to be statistically representative of the universe as a whole.
Remarkably, this breakdown scale can be identified a posteriori using the data. Figure 11 shows that this transition consistently appears at $z\sim 0.01$ across all multipole decompositions of the expansion-rate fluctuation field: the dipole, quadrupole, and octupole all begin to diverge coherently from theoretical predictions below this redshift. As a consequence, the anisotropic expansion-rate structure of the surveyed volume can be effectively described by just seven CC parameters, as soon as the noisy local data are excluded from the analysis.
We have argued that the CC parameters cannot be measured directly from the observer’s position; instead, they must be inferred by fitting the redshift evolution of the expansion rate field over a broad interval. Ensuring that these estimates are independent of the chosen redshift boundaries is therefore essential for achieving genuine convergence to a cosmologically meaningful value. Figures 14 and 15 demonstrate that the best-fit multipoles of the CC parameters, when estimated over redshift intervals with progressively higher lower bounds $z_{\rm min}$, converge toward stable values up to the limit of the data at $z=0.1$.
@@
These results serve as a caution against defining the matter comoving observer in the covariant cosmography framework by averaging data within spherical volumes centered on the Local Group, with radii of approximately $30\,h^{-1}\,\mathrm{Mpc}$ (where, for example, the CF4 sample typically includes nearly 3000 objects.) A more physically motivated estimate of the spatial scale over which one must average to define the velocity of the matter element representing the CC observer can be inferred through a model-dependent argument based on standard cosmological information. We know that the velocity of the matter observer is $v_{o}=188\pm 22$ km/s in the direction $(l,b)=(299^{\circ},5^{\circ})$. Using Cosmicflows-3 sample, [77] determined the contributions to the Local Group (LG) velocity from mass enclosed within a sphere of radius $29h^{-1}$ Mpc, and from the surrounding shell extending from $29h^{-1}$ Mpc to $75h^{-1}$ Mpc (see Table 5). Given the observed CMB dipole, we can deduce that the component of the LG’s motion generated by mass beyond $29h^{-1}$ Mpc is approximately $v_{29}=320\pm 70$ km/s in the direction $(l,b)=(292^{\circ},4^{\circ})$, while that generated by structures beyond $75h^{-1}$ Mpc is about $v_{75}=100\pm 70$ km/s in the direction $(l,b)=(302^{\circ},-38^{\circ})$. The fairly good alignment between these directions and the measured $v_{o}$, and the fact that $v_{75}<v_{o}<v_{29}$ suggests that the effective averaging length corresponding to the size of the fluid element representing the CC observer lies somewhere in the range $29h^{-1}$ Mpc $R~(\mathrm{Mpc})$ $75h^{-1}$ Mpc. This holds under the simplifying assumption that the contribution of shells external to the volume $V$ provides, within statistical noise, a fair representation of the bulk velocity of $V,$ rather than merely the velocity of the Local Group.
@@ 7 Conclusion
Given the large length scale over which the anisotropy signal remains coherent, and the polarity of the multipoles [is?] correlated, it is essential to move beyond traditional linear perturbative analyses framed in the standard model and investigate how the CC parameters relate to the intrinsic parameters of fully relativistic, non-standard metric models that depart from FLRW. We have already begun this exploration in a companion paper [66], where we examine axially symmetric metric models constructed by placing the observer off-center in a Lamaître-Tolman-Bondi (LTB) geometry.